OpenStudy (anonymous):

Work in fractions, and give your answer as a fraction reduced to lowest terms. 1/8+0.6

OpenStudy (anonymous):

.6 = 6/10 1/8 + 6/10 find a common denominator 5/40 + 24/40 = 29/40